# EpiDiff: Enhancing Multi-View Synthesis via Localized Epipolar-Constrained Diffusion

### 11 Dec 2023 v1

@@ Abstract
Generating multiview images from a single view facilitates the rapid generation of a 3D mesh conditioned on a single image. Recent methods [27] that introduce 3D global representation into diffusion models have shown the potential to generate consistent multiviews, but they have reduced generation speed and face challenges in maintaining generalizability and quality. To address this issue, we propose EpiDiff, a localized interactive multiview diffusion model. At the core of the proposed approach is to insert a lightweight epipolar attention block into the frozen diffusion model, leveraging epipolar constraints to enable cross-view interaction among feature maps of neighboring views. The newly initialized 3D modeling module preserves the original feature distribution of the diffusion model, exhibiting compatibility with a variety of base diffusion models. Experiments show that EpiDiff generates 16 multiview images in just 12 seconds, and it surpasses previous methods in quality evaluation metrics, including PSNR, SSIM and LPIPS. Additionally, EpiDiff can generate a more diverse distribution of views, improving the reconstruction quality from generated multiviews. Please see our project page at https://huanngzh.github.io/EpiDiff/.
@@ 1 Introduction
3D reconstruction from a single image has a wide range of applications in fields such as augmented/virtual reality, robotic simulation, and gaming. However, this task presents significant challenges due to the complexity of accurately interpreting spatial structure and details from a 2D perspective observation. It requires extensive knowledge of the 3D world.
Recently, with the introduction of diffusion models [17, 41], the field of 3D generation has experienced rapid and vigorous development. A series of research [30, 34, 68, 11, 28, 1, 5, 7, 12, 15, 21, 33, 35, 58, 20, 67, 2] attempts to train 3D generative diffusion models from scratch on 3D assets. However, due to the limited size of publicly available 3D datasets, these methods demonstrate poor generalizability. Therefore, another group of works [18, 36, 55, 6, 45, 46, 53, 60, 65, 31, 37, 38, 47, 50] propose to distill prior knowledge of 2D diffusion models to create 3D models from text or images via Score Distillation Sampling (SDS) [36]. Although these methods have achieved impressive results, they require expensive per-scene optimization that usually takes tens of minutes or even hours.
@@ 2 Related work
3D generation guided by 2D prior  Given the robust generative capabilities of 2D diffusion models [41, 39, 42], some research efforts directly leverage pretrained image diffusion models. With DreamFusion [36] and SJC [55] proposing to distill a pretrained 2D text-to-image generation model for generating 3D shapes, a series of subsequent studies [23, 6, 45, 46, 53, 60, 65, 31, 37, 38, 47, 50] adopt such a pre-scene optimization pipeline. These methods optimize a 3D representation (e.g., NeRF, mesh, or SDF) conditioned on text or images through Score Distillation Sampling (SDS) [36]. However, the per-scene optimization approach has issues with efficiency. Generating a single scene may take tens of minutes or even hours.
@@ 3.2 Epipolar-constrained attention
Near-Views Cross-Attention  Given a target view, our method relies on the features of the $K$ closest views among the total $N$ views to produce the output feature map. We start with a set of feature maps =$\mathcal{F}=\{F_{i}\mid 1\leq i\leq K\}$, which includes the feature map of the target view $F_{1}$ and those of the $K-1$ nearest reference views. With the camera parameters, we derive rays from patches on the feature map of the target view. We then compute the corresponding epipolar lines on feature maps in $\mathcal{F}$. $S$ points are then sampled uniformly along each epipolar line, such that their re-projections distribute uniformly between the predefined near and far planes. Following this, each sampled point is employed to execute bilinear interpolation on the feature maps of the reference views. This process results in 3D-level features $\mathcal{V}=\{V_{k}^{s}\mid 1\leq k\leq K,1\leq s\leq S\}$. These aligned features, obtained from epipolar geometry projections, are then passed to the nearby views aggregation network. The set $\mathcal{V}$ is split into two parts: $\mathcal{V}_{1}$ with the target view features, and $\mathcal{V}_{2:K}$ with the reference view features. With $\mathcal{V}_{1}$ serving as the query and $\mathcal{V}_{2:K}$ as the key and value, we implement cross-attention on the reshaped feature maps to establish associations, i.e.,
$\widetilde{\mathcal{V}}_{1}=\text{CrossAttention}(\mathcal{V}_{1},\mathcal{V}_{2:K})$ (6)
where $K$ is the number of neighboring views.
@@
Ray Self-Attention  To model spatial points’ contributions to pixel-level features, the ray aggregation module is implemented using self-attention. Upon acquiring the feature representation $\widetilde{\mathcal{V}}_{1}$ of the target view via Near-Views Cross-Attention, we apply self-attention on the features along the ray with dimension of $S$ (the number of points sampled per ray). More formally,
$\overline{\mathcal{V}}_{1}=\text{Attention}(\widetilde{\mathcal{V}}_{1},\widetilde{\mathcal{V}}_{1})$ (7)
This step is crucial for revealing depth relationships among sampled points along rays and capturing their spatial correlations. Finally, we perform a simple weighted summation to fuse the 3D points into 2D feature maps.
@@
Positional Encoding  To augment the module’s capability in comprehending perspective relationships, we incorporate a positional encoding technique derived from Light Field Networks (LFN) [48]. Each pixel is associated with a ray, which is represented using Plücker coordinates. These coordinates are expressed as $r=(o\times d,d)$, where $o$ is the origin of the ray, and $d$ is the direction. The Plücker coordinates are then combined with latent feature maps through a harmonic transformation, enriching the feature representation with spatial information.
@@ 4 Experiments
4.1 Experiment settings
Implementation details  We use Zero123 [25] as our backbone. EpiDiff is trained for 30,000 steps (around 2 days) using 8 80G A800 GPUs, with a learning rate of 1e-5 and a total batch size of 512. In each iteration, we generate 16 multiview images, i.e., $N=16$. Within the ECA Block, the number of sampled rays is aligned with the resolution of feature maps and the number of sampled points per ray $S$ is set to 16. For the Near-Views Cross Attention module, the number of nearby views $K$, is set to 4. For reconstruction, the foreground masks of the generated images are initially predicted using CarveKit. Then, we use the instant-NGP based SDF reconstruction method [14] to reconstruct the shape. Each object undergoes training for 3k steps, which costs about 2 minutes.
4.2 Consistent multiview synthesis
@@
Under the uniform elevation setting, the quantitative results are shown in Tab. 2, and the qualitative results are shown in Fig. 5. Both Zero123 [25] and SyncDreamer [27] struggle to respond to the flexible viewpoint distribution, while our method can still generate relatively realistic and multiview-consistent images. Meanwhile, for the same input image, Our method can generate different plausible instances using different random seeds as shown in Fig. 6.
@@ 4.3 Single view reconstruction
Figure 8: The common failure mode when reconstructing using multiviews at a fixed elevation of $30^{\circ}$ from an input image. When the bottom of the object is round-shaped, the reconstructed bottom appears protruded because the camera’s visible frustum cannot cover the bottom. Using a uniform distribution of elevations can reconstruct a better shape.
@@ 4.4 Ablation study
To assess the effectiveness of our design for ECA Block (Sec. 3.2), we conducted ablation experiments on multiview synthesis at elevation $30^{\circ}$ setting (Tab. 4). We investigated the impact of the number of neighboring views used for aggregating features in Near-Views Cross Attention module, setting $K$ to $\{4,8,16\}$, where $K=16$ indicates the use of all perspective features aggregated to the target view. Experimental results show that increasing $K$ consumes more time and reduces the quality of the generated multiview structure and appearance. Additionally, we examined the positional encoding (PE) strategies and found that removing either the light field PE or the ray-relative PE leads to a decrease in the consistency and quality of the generated images.
@@ 5 Limitations and Conclusion
Conclusion  In this paper, we present EpiDiff, a localized interactive multiview diffusion model designed for generate multiview-consistent and high-quality images from a single-view image. With the Zero123 [25] as the backbone, EpiDiff adopts a lightweight epipolar attention block in the UNet, which utilizes epipolar contraints to enable cross-view interaction among feature maps of neighboring views. This module models 3D consistency whinin the original feature space of the diffusion model, exhitbiting adaptability to various base diffusion models. Extensive experiments demonstrate that EpiDiff not only efficiently generates consistent and high-quality multiview images, but is also capable of generating more freely distributed views, help to better reconstruction from generated multiviews.

### 16 Mar 2024 v2

@@ Abstract
Generating multiview images from a single view facilitates the rapid generation of a 3D mesh conditioned on a single image. Recent methods [27] that introduce 3D global representation into diffusion models have shown the potential to generate consistent multiviews, but they have reduced generation speed and face challenges in maintaining generalizability and quality. To address this issue, we propose EpiDiff, a localized interactive multiview diffusion model. At the core of the proposed approach is to insert a lightweight epipolar attention block into the frozen diffusion model, leveraging epipolar constraints to enable cross-view interaction among feature maps of neighboring views. The newly initialized 3D modeling module preserves the original feature distribution of the diffusion model, exhibiting compatibility with a variety of base diffusion models. Experiments show that EpiDiff generates 16 multiview images in just 12 seconds, and it surpasses previous methods in quality evaluation metrics, including PSNR, SSIM and LPIPS. Additionally, EpiDiff can generate a more diverse distribution of views, improving the reconstruction quality from generated multiviews. Please see the project page at huanngzh.github.io/EpiDiff/.
@@ 1 Introduction
3D reconstruction from a single image has a wide range of applications in augmented/virtual reality, robotic simulation, and gaming. However, this task presents significant challenges due to the complexity of accurately interpreting spatial structure and details from a 2D perspective observation. It requires extensive knowledge of the 3D world.
Recently, with the introduction of diffusion models [17, 41], the field of 3D generation has experienced rapid and vigorous development. A series of research [30, 34, 68, 11, 28, 1, 5, 7, 12, 15, 21, 33, 35, 58, 20, 67, 2] attempts to train 3D generative diffusion models from scratch on 3D assets. However, due to the limited size of publicly available 3D datasets, these methods demonstrate poor generalizability. Therefore, another group of works [18, 36, 55, 6, 60, 31, 37, 38, 50, yu2023csd, sun2023dreamcraft3d] propose to distill prior knowledge of 2D diffusion models to create 3D models from text or images via Score Distillation Sampling (SDS) [36]. Although these methods have achieved impressive results, they require expensive per-scene optimization that usually takes tens of minutes or even hours.
@@ 2 Related work
3D generation guided by 2D prior  Given the robust generative capabilities of 2D diffusion models [41, 39, 42], some research efforts directly leverage pretrained image diffusion models. With DreamFusion [36] and SJC [55] proposing to distill a pretrained 2D text-to-image generation model for generating 3D shapes, a series of subsequent studies [23, 6, 60, 31, 37, 38, 50, yu2023csd, sun2023dreamcraft3d] adopt such a pre-scene optimization pipeline. These methods optimize a 3D representation (e.g., NeRF, mesh, or SDF) conditioned on text or images through Score Distillation Sampling (SDS) [36]. However, the per-scene optimization approach has issues with efficiency. Generating a single scene may take tens of minutes or even hours.
@@ 3.2 Epipolar-constrained attention
Near-Views Cross-Attention  Given a target view, our method relies on the features of the $F$ closest views among the total $N$ views to produce the output feature map. We start by emitting rays from patches on the target view, and sample $S$ points uniformly along each ray. We then project them onto $F-1$ neighboring views. For each point $s$, we get the feature $v^{s}$ by positional encoding and projected feature $\{f^{s}_{i}\mid 1\leq i\leq F-1\}$ on reference views. We enable cross-view interaction by a cross-attention layer. Specifically, using point feature $v^{s}$ as query, we implement $\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(\frac{QK^{T}}{\sqrt{d}})\cdot V$ with
$Q=W^{Q}v^{s},K=W^{K}[f^{s}_{1},\cdot\cdot,f^{s}_{F-1}],V=W^{V}[f^{s}_{1},\cdot\cdot,f^{s}_{F-1}],$ (6)
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation operation, and $W^{Q}$, $W^{K}$, $W^{V}$ are learnable matrices that project the inputs to query, key, and value, respectively.
@@
Ray Self-Attention  To model spatial points’ contributions to pixel-level features, the ray aggregation module is implemented using self-attention. Upon acquiring the feature of each sampled point $\widetilde{v}^{s}(1\leq s\leq S)$, we apply self-attention along the ray with dimension of $S$, deriving all of query, key and value features from $[\widetilde{v}^{1},\cdots,\widetilde{v}^{S}]$. This step is crucial for revealing depth relationships among sampled points along rays and capturing their spatial correlations. Finally, we perform a weighted summation to fuse 3D features into 2D feature of target patch $\widetilde{f}$.
@@
Positional Encoding  To enhance the module’s understanding of perspective relationships, we have adopted a positional encoding technique inspired by Light Field Networks (LFN) [48]. This approach is distinct in that each pixel is associated with a ray, utilizing Plücker coordinates for representation, denoted as $r=(o\times d,d)$, where $o$ signifies the ray’s origin, and $d$ its direction. Unlike traditional position encoding methods where all pixels of an image share the same positional encoding, the LFN [48] method employs MVP transformation to back-project each pixel of the image into the world coordinate system, precisely locating the corresponding ray for each pixel. It introduces strict geometric constraints and sufficiently incorporates camera parameters, effectively applying encoding at the pixel level. By integrating these encoded coordinates obtained by LFN with latent feature maps via a harmonic transformation, our method significantly enriches the spatial information within the feature representation.
@@ 4 Experiments
4.1 Experiment settings
Implementation details  We use Zero123 [25] as our backbone. EpiDiff is trained for 30,000 steps (around 2 days) using 8 80G A800 GPUs, with a learning rate of 1e-5 and a total batch size of 512. In each iteration, we generate 16 multiview images, i.e., $N=16$. Within the ECA Block, the number of sampled rays is aligned with the resolution of feature maps and the number of sampled points per ray $S$ is set to 16. For the Near-Views Cross Attention module, the number of nearby views $F$, is set to 4. For reconstruction, the foreground masks of the generated images are initially predicted using CarveKit. Then, we use the instant-NGP based SDF reconstruction method [14] to reconstruct the shape. Each object undergoes training for 3k steps, which costs about 2 minutes.
4.2 Consistent multiview synthesis
@@
Under the uniform elevation setting, the quantitative results are shown in Tab. 2, and the qualitative results are shown in Fig. 5. The uniform view setting offers a more flexible viewpoint distribution. SyncDreamer [27] lacks precise cross-view geometric constraints, leading to overfitting training viewpoints and a bias towards top-down perspectives in multiview generation, as shown in Fig. 5. Our method can still generate relatively realistic and multiview-consistent images, benefiting from (a) the introduction of localized epipolar geometric constraints along with effective camera position encoding, and (b) the random viewpoint sampling training method. Meanwhile, for the same input image, Our method can generate different plausible instances using different random seeds as shown in Fig. 6.
@@ 4.4 Ablation study
To assess the effectiveness of our design for ECA Block (Sec. 3.2), we conducted ablation experiments on multiview synthesis at elevation $30^{\circ}$ setting (Tab. 4). We investigated the impact of the number of neighboring views used for aggregating features in Near-Views Cross Attention module, setting $F$ to $\{4,8,16\}$, where $K=16$ indicates the use of all perspective features aggregated to the target view. Experimental results show that increasing $F$ consumes more time and reduces the quality of the generated multiview structure and appearance. Additionally, we examined the positional encoding (PE) strategies and found that removing either the light field PE or the ray-relative PE leads to a decrease in the consistency and quality of the generated images.
@@ 5 Limitations and Conclusion
Figure 8: The common failure mode when reconstructing using multiviews at a fixed elevation of $30^{\circ}$. When the bottom of the object is round-shaped, the reconstructed bottom appears protruded because the camera’s visible frustum can not cover the bottom. Using a uniform distribution of elevations can solve it.
Conclusion  In this paper, we present EpiDiff, a localized interactive multiview diffusion model designed for generating multiview-consistent and high-quality images from a single-view image. With the Zero123 [25] as the backbone, EpiDiff adopts a lightweight epipolar attention block in the UNet, which utilizes epipolar contraints to enable cross-view interaction among feature maps of neighboring views. This module models 3D consistency whinin the original feature space of the diffusion model, exhitbiting adaptability to various base diffusion models. Extensive experiments demonstrate that EpiDiff not only efficiently generates consistent multiview images, but is also capable of generating more freely distributed views, help to better reconstruction from generated multiviews.